A reactant decomposes with a half-life of 39.5

39.5 s when its initial concentration is 0.322

0.322 M. When the initial concentration is 0.790

0.790 M, this same reactant decomposes with a half‑life of 96.9

96.9 s.

What is the order of the reaction?

1

2


What is the value and unit of the rate constant for this reaction?


1
Expert's answer
2021-04-19T04:34:37-0400

From given data .


Half life is directly Proportional to Initial Concentration .


So, Order is " 2 " .


For Second order .

Unit of rate Constant = mol-1 . lit1 . Sec-1 .


K = 1 / initial Concentration × Half life .


K = 1 / 0.322 × 39.5


K= 1 / 12.719


K= 0.079 mol- . lit .sec-1 .